# Curious QNEIs from QNEC: New Bounds on Null Energy in Quantum Field Theory

### v1

@@ 1 Introduction
Energy plays a ubiquitous role in physics; it underlies dynamics in all systems, connects matter to geometry in gravity, and is central to thermodynamics and statistical physics. Many physical classical field theories obey pointwise energy conditions, and these have played an important role in, for example, singularity theorems penrose1965gravitational . However, for local, relativistic quantum field theories (QFTs), the study of energy is both richer and more precarious. Due to the short-distance entanglement of the vacuum, the energy-momentum density at any point can be made arbitrarily negative epstein1965nonpositivity .
To get a lower bound on the stress tensor operator in QFTs, it must be ‘smeared’ – integrated against a test function – over a spacetime domain. Such constraints are called quantum energy inequalities (QEIs) Fewster:2012yh . In general, a QEI can express the lower bound on the energy density in terms of the expectation values of other operators; however, particularly powerful are state-independent QEIs which express smeared energy densities in terms of a state-independent functional of the smearing, schematically:
@@
where $v^{\mu}$ is the null vector generating the geodesic. Among other applications, achronal ANEC is necessary for causality in AdS/CFT Gao:2000ga , and forbids long traversable wormholes morris1988wormholes ; Hochberg:1998ii . However, of additional interest are QEIs of the more semi-local variety, where, unlike in ANEC, $D$ is a compact domain. Due to their state-independence and semi-locality, such QEIs have been a key tool in extending the classical singularity theorems of general relativity (such as those by Hawking and Penrose penrose1965gravitational ; Hawking:1966sx ; Hawking:1966jv ; Hawking:1967ju ) to semi-classical gravity coupled to QFT Brown:2018hym ; Fewster:2019bjg ; Fewster:2021mmz .
While the ANEC has been shown to hold for all QFTs in Minkowski spacetime Faulkner:2016mzt ; Hartman:2016lgu , and for free theories in certain classes of curved spacetimes Wald:1991xn ; Kelly:2014mra ; Kontou:2015yha , deriving and proving QEIs of the semi-local variety is a difficult endeavour. Perhaps the best known example is Fewster and Hollands’ proof that all two-dimensional conformal field theories (CFTs) admit the following semi-local QEI on the stress tensor Fewster:2004nj
@@
where $c$ is the central charge of the CFT. See also Blanco:2017akw for a stronger but state-dependent version of (3), and Blanco:2013lea , which derives bounds on how far negative energy can be isolated away from positive energy. Both of those papers use monotonicity of relative entropy, which we will use too. The first result in this paper will be a new, elementary proof and extension of (3).
A different extension of the FH bound to states of $d>2$ QFTs consistently coupled to semi-classical gravity, coined the smeared null energy condition (SNEC), was conjectured by Freivogel and Krommydas Freivogel:2018gxj
@@
in which the central charge is replaced by the inverse Newton’s constant. The key differences with respect to the FH bound are the coupling to gravity and that $d>2$. The SNEC has been proven for holographic CFTs coupled to induced gravity in a braneworld model Leichenauer:2018tnq , however, broader evidence for the SNEC has so far remained elusive. This is partially because, as a gravitational bound, gravitational physics must be part of its derivation.
Amongst the QEIs that can be established purely in QFTs in greater than two dimensions, almost all results have been established for free QFTs (with ANEC the notable counterexample). Furthermore, it is known that certain free theories do not admit state-independent lower bounds: an example is the non-minimally coupled scalar field for which one can construct states of arbitrarily negative expectation value of the timelike or null-contracted stress tensor, regardless of the domain over which it is smeared Fewster:2007ec ; Fliss:2023rzi . This non-existence of a QEI is pertinent to the free scalar CFT even in flat space, where the non-minimal coupling manifests as an improvement term needed to make the stress tensor traceless.
It is natural to wonder whether this QEI non-existence is particular to free field theories, or whether QEIs are generically too much to hope for. Before this paper, this question remained largely unanswered, though investigations into large-$N$ CFTs with a ‘generalized free field’ structure have indicated that interactions may prevent the ‘pile-up’ of negative energy density allowed in free theories Fliss:2024dxe .
We will derive new QEIs on null energy, and so answer this question. To do so, we will make use of a completely complementary approach from those used so far in deriving QEIs, one that makes use of the connection between energy and entropy, through the quantum null energy condition (QNEC) Bousso:2015mna , which bounds the local expectation value of the null stress tensor by a second-order shape derivative of the entanglement entropy at that point:
@@
The above formula is schematic; in the body of this paper, we will provide more precise details. The QNEC lends a universality to our results, because it has been proven for free and super-renormalisable Bousso:2015wca , holographic Koeller:2015qmn , and generic interacting QFTs Balakrishnan:2017bjg ; Ceyhan:2018zfg ; Hollands:2025glm . Moreover, for $d>2$ interacting theories, there is strong evidence that this local form of the QNEC is saturated Leichenauer:2018obf ; Balakrishnan:2019gxl , implying that it is the tightest possible bound on local energy densities. Note that ANEC follows from integrating QNEC. The entanglement entropy that appears in (5) obeys its own inequalities – weak monotonicity, strong subadditivity (SSA), and others – and we will leverage these in our derivations to go from entropy-dependent to state-independent bounds.
In the gravitational context, QNEC, and the quantum focussing conjecture (QFC) from which it descends Bousso:2015mna , connect energy and entropy, which has been insightful for holography and semi-classical gravity Faulkner:2022mlp . Related inequalities, such as the generalised second law Bekenstein:1974ax and the Bousso bound Bousso:1999xy ; Rolph:2022csa , are also effective starting points for constraining semi-classical gravity in a manner complementary to QEI-based approaches Wall:2010jtc ; Bousso:2022cun ; Bousso:2022tdb ; Bousso:2025xyc .
However, from a purely field-theoretic, non-gravitational point of view, it is not clear that the QNEC, even when it is saturated, is an effective bound on the energy density. For one, there are states for which the right-hand side of (5) is arbitrarily negative (consistent with the pointwise stress tensor being an unbounded operator). Secondly, at a practical level, the evaluation of entanglement entropy is extremely difficult, except for certain symmetric regions and for holographic theories Ryu:2006bv . Lastly, there is an undesirable mismatch between the linear (on the left-hand side) and non-linear (on the right-hand side) state dependences of (5).
@@
Firstly, in Section 2.1, we show that the ‘conformally improved’ QNEC Wall:2011kb directly implies the FH bound (3) and extends it to all super-renormalizable QFTs. Our derivation is extremely simple, yet, as far as we are aware, has not yet been presented in the literature.
@@
$\boxed{\int\text{d}v\,m_{[g,\zeta]}(v)\langle T_{vv}\rangle\geq-\frac{c_{\scriptscriptstyle\mathrm{UV}}}{12\pi}\int\text{d}v\frac{g^{\prime}(v)}{v-\zeta(v)}~,}$ (6)
where $c_{\scriptscriptstyle\mathrm{UV}}$ is the central charge of the UV-fixed point of the theory and $m_{[g,\zeta]}(v)$ is a convolution of $g$ with a kernel associated to the modular Hamiltonian of a null strip whose precise form is given in (59). Much like $g$ itself, $m$ is positive and localised (e.g. has compact support when $g$ has compact support). The function $\zeta$ essentially parametrizes the choice of null strips as entangling regions in our monotonicity argument and appears in the definition of $m$.
@@
In Section 3, we adapt our methods to interacting CFTs in $d>2$ dimensions. By considering the monotonicity argument applied to ‘nearly null’ strips, we derive an infinite family of QNEIs smeared over a causal diamond uniform in d2d-2 transverse directions. Again, to be specific, given a positive and localized function $g(u,v)$ of the lightcone coordinates, we derive
$\boxed{\begin{split}\int\text{d}^{d-2}y_{\perp}&\int\text{d}u\text{d}v\mathcal{M}_{[g,\zeta]}(u,v)\langle T_{vv}\rangle\\ &\geq-\frac{\beta\,c_{\scriptscriptstyle\mathrm{T}}}{2\pi}\int\text{d}^{d-2}y_{\perp}\int\text{d}u\text{d}v\left(\frac{\partial_{v}g(u,v)}{\epsilon_{u}(u,v)^{\frac{d-2}{2}}(v-\zeta(u,v))^{\frac{d}{2}}}+O(\epsilon_{u})\right)~,\end{split}}$ (7)
where $c_{\scriptscriptstyle\mathrm{T}}$ is the stress tensor two-point coefficient, and $\beta$ is a theory-dependent constant that we will define later. $\mathcal{M}$ is defined completely analogously as $m$ in our $d=2$ bound: it is a particular convolution of $g$ with a kernel associated to the modular Hamiltonian of a null strip and depends on the choice of $\zeta$. Its precise definition is given in (3.2). The new ingredient in (7) in comparison to (6) is the additional function $\epsilon_{u}$, and that (7) applies when $\epsilon_{u}\ll 1$. The origin of $\epsilon_{u}$ is a parameterization of how far away the ‘nearly null’ strip regions in our derivation are from being truly null. We discuss relations of this new family of higher-dimensional QNEIs to the recently proposed double smeared energy condition (DSNEC) Fliss:2021gdz ; Fliss:2021phs . While the uniform transverse smearing imposes some limitations on the utility of our results (essentially introducing a divergent IR volume to the bound), to our knowledge, our results present both the first instances of QEIs derived taking the QNEC as an input, and the first QEIs in a large class of interacting theories in greater than two dimensions. We see this as a major step in addressing the existence of state-independent bounds on locally smeared energy density in generic interacting theories.
@@ 2.1 Fewster-Hollands bound from 2d QNEC
Consider a QFT on 2d Minkowski spacetime and let $T_{vv}(v)$ be the $vv$ component of its stress tensor at the point $p=(0,v)$. The choice of $u=0$ is without loss of generality. The QNEC for 2d QFTs is Wall:2011kb ; Bousso:2015mna ; Koeller:2015qmn
@@
The 2d QNEC (12) has been proven for holographic 2d CFTs, and relevant deformations thereof, in Koeller:2015qmn , but not in complete generality. Compared to QNEC in higher dimensions, the 2d QNEC is special because it has the $(\partial_{v}S)^{2}$ term, which is absent from the higher-dimensional QNEC. In CFTs, the 2d QNEC is saturated by Bañados states (Virasoro coadjoint orbit states) Khandker:2018xls ; Ecker:2019ocp . The $(\partial_{v}S)^{2}$ term makes the right-hand side of (12) transform as a conformal primary, which means that both sides of the inequality have the same conformal anomaly, making it independent of the conformal frame Wall:2011kb .
@@
This is precisely the Fewster-Hollands (FH) bound for 2d CFTs Fewster:2004nj introduced in Section 1.
@@
  • Unlike the 2d QNEC,  (12), the FH bound is not saturated for the vacuum state for the simple reason that $\langle T_{vv}\rangle_{\Omega}=0$. However at the UV fixed point, the bound is tight and the states saturating it can easily be constructed jrf_modave as
@@
  • The FH bound (14) only holds in two dimensions; there are explicit counterexamples in higher dimensions. Namely, it was shown by Fewster and Roman that the left-hand side of (14) is unbounded in the class of “0+2 states” (superpositions of zero and two-particle states) of free scalar field theory in $d=4$ dimensions Fewster:2002ne . More generally, one can construct a series of squeezed states in free scalar field theories of arbitrary dimension $d>2$ where the left-hand side of (14) can be made arbitrarily negative Fliss:2021gdz . However, when the free theory comes equipped with a UV cutoff on the transverse momentum, then the properties of free theories quantized on lightsheets Burkardt:1995ct ; Wall:2011hj allow one to “bootstrap” the FH bound to higher dimensions:
@@
  • Our derivation of (14) is novel because it is both elementary and extends the proof of this FH bound to all 2d QFTs for which (12) is proven, which includes some non-conformal theories, namely relevant deformations of holographic 2d CFTs Koeller:2015qmn .
  • Lastly, as mentioned above, the coefficient $c_{\scriptscriptstyle\mathrm{UV}}$ appearing in (14) is the central charge of the QFT’s UV fixed point. However, at the other end of the RG flow, the QFT is described by an IR CFT with central charge $c_{\scriptscriptstyle\mathrm{IR}}$. The stress tensor of this theory will satisfy the FH bound with coefficient $c_{\scriptscriptstyle\mathrm{IR}}$, which can be proven directly in the CFT. This is consistent with our claim due to the monotonicity of the central charge under RG flows Zamolodchikov:1986gt ,
@@ 2.2 An family of 2d QNEIs
In the interest of investigating the utility of the QNEC for establishing QNEIs in higher dimensions, we will look at deriving a version of the FH bound that does not make use of the $(\partial_{v}S)^{2}$ term of (12) since this term doesn’t appear in $d>2$. If it did, then, using the same method as in section 2.1, we would have a state-independent lower bound on $\int dvg(v)\langle T_{vv}\rangle$ in $d>2$, which is impossible Fewster:2002ne . We will work exclusively in $d=2$ QFTs in this subsection, leaving an exploration to higher-dimensional QFTs in Section 3. We will find that this will lead to an infinite family of QEIs, parameterized by a function $\zeta$, which are interesting in their own right.
@@
So, with respect to the FH 2d bound, the QNEIs we derive will be proven for any 2d QFT for which (20) is proven, which includes super-renormalisable QFTs Bousso:2014sda .

### v2

@@ 1 Introduction
Energy plays a ubiquitous role in physics; it underlies dynamics in all systems, connects matter to geometry in gravity, and is central to thermodynamics and statistical physics. Many physical classical field theories obey pointwise energy conditions, and these have played an important role in, for example, singularity theorems Penrose (1965). However, for local, relativistic quantum field theories (QFTs), the study of energy is both richer and more precarious. Due to the short-distance entanglement of the vacuum, the energy-momentum density at any point can be made arbitrarily negative Epstein et al. (1965).
To get a lower bound on the stress tensor operator in QFTs, it must be ‘smeared’ – integrated against a test function – over a spacetime domain. Such constraints are called quantum energy inequalities (QEIs) Fewster (2012). In general, a QEI can express the lower bound on the energy density in terms of the expectation values of other operators; however, particularly powerful are state-independent QEIs which express smeared energy densities in terms of a state-independent functional of the smearing, schematically:
@@
where $v^{\mu}$ is the null vector generating the geodesic. Among other applications, achronal ANEC is necessary for causality in AdS/CFT Gao and Wald (2000), and forbids long traversable wormholes Morris et al. (1988); Hochberg and Visser (1998). However, of additional interest are QEIs of the more semi-local variety, where, unlike in ANEC, $D$ is a compact domain. Due to their state-independence and semi-locality, such QEIs have been a key tool in extending the classical singularity theorems of general relativity (such as those by Hawking and Penrose Penrose (1965); Hawking (1966b, a, 1967)) to semi-classical gravity coupled to QFT Brown et al. (2018); Fewster and Kontou (2020, 2022).
While the ANEC has been shown to hold for all QFTs in Minkowski spacetime Faulkner et al. (2016a); Hartman et al. (2017), and for free theories in certain classes of curved spacetimes Wald and Yurtsever (1991); Kelly and Wall (2014); Kontou and Olum (2015), deriving and proving QEIs of the semi-local variety is a difficult endeavour. Perhaps the best known example is Fewster and Hollands’ proof that all two-dimensional conformal field theories (CFTs) admit the following semi-local QEI on the stress tensor Fewster and Hollands (2005)
@@
where $c$ is the central charge of the CFT. See also Blanco et al. (2018) for a stronger but state-dependent version of (3), and Blanco and Casini (2013), which derives bounds on how far negative energy can be isolated away from positive energy. Both of those papers use monotonicity of relative entropy, which we will use too. The first result in this paper will be a new, elementary proof and extension of (3).
A different extension of the FH bound to states of $d>2$ QFTs consistently coupled to semi-classical gravity, coined the smeared null energy condition (SNEC), was conjectured by Freivogel and Krommydas Freivogel and Krommydas (2018)
@@
in which the central charge is replaced by the inverse Newton’s constant. The key differences with respect to the FH bound are the coupling to gravity and that $d>2$. The SNEC has been proven for holographic CFTs coupled to induced gravity in a braneworld model Leichenauer and Levine (2019), however, broader evidence for the SNEC has so far remained elusive. This is partially because, as a gravitational bound, gravitational physics must be part of its derivation.
Amongst the QEIs that can be established purely in QFTs in greater than two dimensions, almost all results have been established for free QFTs (with ANEC the notable counterexample). Furthermore, it is known that certain free theories do not admit state-independent lower bounds: an example is the non-minimally coupled scalar field for which one can construct states of arbitrarily negative expectation value of the timelike or null-contracted stress tensor, regardless of the domain over which it is smeared Fewster and Osterbrink (2008); Fliss et al. (2024b). This non-existence of a QEI is pertinent to the free scalar CFT even in flat space, where the non-minimal coupling manifests as an improvement term needed to make the stress tensor traceless.
It is natural to wonder whether this QEI non-existence is particular to free field theories, or whether QEIs are generically too much to hope for. Before this paper, this question remained largely unanswered, though investigations into large-$N$ CFTs with a ‘generalized free field’ structure have indicated that interactions may prevent the ‘pile-up’ of negative energy density allowed in free theories Fliss et al. (2024a).
We will derive new QEIs on null energy, and so answer this question. To do so, we will make use of a completely complementary approach from those used so far in deriving QEIs, one that makes use of the connection between energy and entropy, through the quantum null energy condition (QNEC) Bousso et al. (2016b), which bounds the local expectation value of the null stress tensor by a second-order shape derivative of the entanglement entropy at that point:
@@
The above formula is schematic; in the body of this paper, we will provide more precise details. The QNEC lends a universality to our results, because it has been proven for free and super-renormalisable Bousso et al. (2016a), holographic Koeller and Leichenauer (2016), and generic interacting QFTs Balakrishnan et al. (2019); Ceyhan and Faulkner (2020); Hollands and Longo (2025). Moreover, for $d>2$ interacting theories, there is strong evidence that this local form of the QNEC is saturated Leichenauer et al. (2018); Balakrishnan et al. (2022), implying that it is the tightest possible bound on local energy densities. Note that ANEC follows from integrating QNEC. The entanglement entropy that appears in (5) obeys its own inequalities – weak monotonicity, strong subadditivity (SSA), and others – and we will leverage these in our derivations to go from entropy-dependent to state-independent bounds.
In the gravitational context, QNEC, and the quantum focussing conjecture (QFC) from which it descends Bousso et al. (2016b), connect energy and entropy, which has been insightful for holography and semi-classical gravity Faulkner et al. (2022). Related inequalities, such as the generalised second law Bekenstein (1974) and the Bousso bound Bousso (1999); Rolph (2022), are also effective starting points for constraining semi-classical gravity in a manner complementary to QEI-based approaches Wall (2013); Bousso and Shahbazi-Moghaddam (2022, 2023); Bousso (2025).
However, from a purely field-theoretic, non-gravitational point of view, it is not clear that the QNEC, even when it is saturated, is an effective bound on the energy density. For one, there are states for which the right-hand side of (5) is arbitrarily negative (consistent with the pointwise stress tensor being an unbounded operator). Secondly, at a practical level, the evaluation of entanglement entropy is extremely difficult, except for certain symmetric regions and for holographic theories Ryu and Takayanagi (2006). Lastly, there is an undesirable mismatch between the linear (on the left-hand side) and non-linear (on the right-hand side) state dependences of (5).
@@
Firstly, in Section 2.1, we show that the ‘conformally improved’ QNEC Wall (2012b) directly implies the FH bound (3) and extends it to all super-renormalizable QFTs. Our derivation is extremely simple, yet, as far as we are aware, has not yet been presented in the literature.
@@
$\boxed{\int\text{d}v\,m(v)\langle T_{vv}\rangle\geq-\frac{c_{\scriptscriptstyle\mathrm{UV}}}{12\pi}\int\text{d}v\frac{g^{\prime}(v)}{v-\zeta(v)}~,}$ (6)
where $c_{\scriptscriptstyle\mathrm{UV}}$ is the central charge of the UV-fixed point of the theory and $m(v)$ is a convolution of $g$ with a kernel associated to the modular Hamiltonian of a null strip whose precise form is given in (59). Much like $g$ itself, $m$ is positive and localised (e.g. has compact support when $g$ has compact support). The function $\zeta$ essentially parametrizes the choice of null strips as entangling regions in our monotonicity argument and appears in the definition of $m$.
@@
In Section 3, we adapt our methods to interacting CFTs in $d>2$ dimensions. By considering the monotonicity argument applied to ‘nearly null’ strips, we derive an infinite family of QNEIs smeared over a causal diamond uniform in d2d-2 transverse directions. Again, to be specific, given a positive and localized function $g(v)$ of the lightcone coordinates, we derive
$\boxed{\begin{split}\int\text{d}^{d-2}y_{\perp}&\int\text{d}v\mathcal{M}(v)\langle T_{vv}\rangle\\ &\geq-\frac{\beta\,c_{\scriptscriptstyle\mathrm{T}}}{2\pi}\int\text{d}^{d-2}y_{\perp}\int\text{d}v\left(\frac{g^{\prime}(v)}{\epsilon_{u}(v)^{\frac{d-2}{2}}(v-\zeta(v))^{\frac{d}{2}}}+O(\epsilon_{u})\right)~,\end{split}}$ (7)
where $c_{\scriptscriptstyle\mathrm{T}}$ is the stress tensor two-point coefficient, and $\beta$ is a theory-dependent constant that we will define later. $\mathcal{M}$ is defined completely analogously as $m$ in our $d=2$ bound: it is a particular convolution of $g$ with a kernel associated to the modular Hamiltonian of a null strip and depends on the choice of $\zeta$. Its precise definition is given in (3.2). The new ingredient in (7) in comparison to (6) is the additional function $\epsilon_{u}$, and that (7) applies when $\epsilon_{u}\ll 1$. The origin of $\epsilon_{u}$ is a parameterization of how far away the ‘nearly null’ strip regions (that appear at an intermediate step in our derivation) are from being truly null. The leading-order, non-positive-power terms in (7) are zero or state-independent, but the linear order $\epsilon_{u}$ terms are state-dependent; our bound is approximately state-independent for small $\epsilon_{u}$. While the uniform transverse smearing imposes some limitations on the utility of our results (essentially introducing a divergent IR volume to the bound), to our knowledge, our results present both the first instances of QEIs derived taking the QNEC as an input, and the first QEIs in a large class of interacting theories in greater than two dimensions. We see this as a major step in addressing the existence of state-independent bounds on locally smeared energy density in generic interacting theories.
@@ 2.1 Fewster-Hollands bound from 2d QNEC
Consider a QFT on 2d Minkowski spacetime and let $T_{vv}(v)$ be the $vv$ component of its stress tensor at the point $p=(0,v)$. The choice of $u=0$ is without loss of generality. The QNEC for 2d QFTs is Wall (2012b); Bousso et al. (2016b); Koeller and Leichenauer (2016)
@@
The 2d QNEC (12) has been proven for 2d CFTs and relevant deformations of holographic 2d CFTs Koeller and Leichenauer (2016), but not in complete generality. Compared to QNEC in higher dimensions, the 2d QNEC is special because it has the $(\partial_{v}S)^{2}$ term, which is absent from the higher-dimensional QNEC. In CFTs, the 2d QNEC is saturated by Bañados states (Virasoro coadjoint orbit states) Khandker et al. (2018); Ecker et al. (2019). The $(\partial_{v}S)^{2}$ term makes the right-hand side of (12) transform as a conformal primary, which means that both sides of the inequality have the same conformal anomaly, making it independent of the conformal frame Wall (2012b).
@@
This is precisely the Fewster-Hollands (FH) bound for 2d CFTs Fewster and Hollands (2005) introduced in Section 1.
@@
  • Unlike the 2d QNEC,  (12), the FH bound is not saturated for the vacuum state for the simple reason that $\langle T_{vv}\rangle_{\Omega}=0$. However, at the UV fixed point, the bound is tight and the states saturating it can easily be constructed Fliss as
@@
  • The FH bound (14) only holds in two dimensions; there are explicit counterexamples in higher dimensions. Namely, it was shown by Fewster and Roman that the left-hand side of (14) is unbounded in the class of “0+2 states” (superpositions of zero and two-particle states) of free scalar field theory in $d=4$ dimensions Fewster and Roman (2003). More generally, one can construct a series of squeezed states in free scalar field theories of arbitrary dimension $d>2$ where the left-hand side of (14) can be made arbitrarily negative Fliss and Freivogel (2022). However, when the free theory comes equipped with a UV cutoff on the transverse momentum, then the properties of free theories quantized on lightsheets Burkardt (1996); Wall (2012a) allow one to “bootstrap” the FH bound to higher dimensions:
@@
  • Our derivation of (14) is novel because it is both elementary and extends the proof of this FH bound to all 2d QFTs for which (12) is proven, which includes some non-conformal theories, namely relevant deformations of holographic 2d CFTs Koeller and Leichenauer (2016).
  • Lastly, as mentioned above, the coefficient $c_{\scriptscriptstyle\mathrm{UV}}$ appearing in (14) is the central charge of the QFT’s UV fixed point. However, at the other end of the RG flow, the QFT is described by an IR CFT with central charge $c_{\scriptscriptstyle\mathrm{IR}}$. The stress tensor of this theory will satisfy the FH bound with coefficient $c_{\scriptscriptstyle\mathrm{IR}}$, which can be proven directly in the CFT. This is consistent with our claim due to the monotonicity of the central charge under RG flows Zamolodchikov (1986),
@@ 2.2 An family of 2d QNEIs
In the interest of investigating the utility of the QNEC for establishing QNEIs in higher dimensions, we will look at deriving a version of the FH bound that does not make use of the $(\partial_{v}S)^{2}$ term of (12) since this term doesn’t appear in $d>2$. If it did, then, using the same method as in section 2.1, we would have a state-independent lower bound on $\int dvg(v)\langle T_{vv}\rangle$ in $d>2$, which is impossible Fewster and Roman (2003). We will work exclusively in $d=2$ QFTs in this subsection, leaving an exploration to higher-dimensional QFTs in Section 3. We will find that this will lead to an infinite family of QEIs, parameterized by a function $\zeta$, which are interesting in their own right.
@@
So, with respect to the FH 2d bound, the QNEIs we derive will be proven for any 2d QFT for which (20) is proven, which includes super-renormalisable QFTs Bousso et al. (2014).